# Many-Body Localization in the Age of Classical Computing

### v2

@@ VI.1.2 Thouless time
Figure 5: Spectral form factor $K(\tau)$ in the $J_{1}$-$J_{2}$ spin chain. (a) Deep in the ergodic regime, for $W=2.5$, the SFF of the $J_{1}$-$J_{2}$ model has the ramp-plateau structure following the GOE prediction (31) (denoted by gray dashed line) for $\tau>t_{Th}/t_{H}$, defining the Thouless time $t_{Th}$. The points $\tau_{Th}=t_{Th}/t_{H}$ are denoted by red circles, while the red square denotes $\tau_{\mathrm{H}}=1$. (b) Deep in the MBL regime of $J_{1}$-$J_{2}$ model, for $W=10$, the SFF is slowly tending, with an increase of $L$, to the result $K_{PS}(\tau)=1$ for Poisson level statistics. Convergence of SFF to $K_{PS}(\tau)$ for $L\to\infty$ for $W>W_{c}$ would imply the presence of the MBL phase in the system.
@@
A spectral form factor (SFF) is a Fourier transform of a two-point spectral correlation function. It quantifies long-range correlations between eigenvalues of many-body systems, and it provides insight into the dynamical properties at timescales below the Heisenberg time $t_{H}$. Investigations of the SFF played a significant role in our understanding of the ETH-MBL crossover. If we denote the unfolded [^16] Spectral unfolding is a procedure of setting the density of eigenvalues to unity [258, 237]. eigenvalues of the considered quantum system by $\{\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{\mathcal{N}}\}$, the averaged SFF is defined as
@@
where the average is taken over disorder realizations, $g(\epsilon)$ is a Gaussian function reducing the influence of the spectrum’s edges, and $Z$ is a normalization factor ensuring that $K(\tau)\stackrel{{\scriptstyle\tau\to\infty}}{{\longrightarrow}}1$. The SFF is known analytically for the Gaussian ensembles of random matrices. In particular, for the GOE, it reads [410]:
@@
The Thouless time is usually understood as the longest physically relevant relaxation time in the system. Its notion was initially introduced in the context of Anderson localization in non-interacting models as a time scale at which a particle reaches the boundary of the system [185, 655]. The 3D Anderson model analysis shows that the Thouless time, extracted from the SFF, follows the expected scalings. Consistently with the presence of diffusion in the delocalized phase of the 3D Anderson model, [593, 738] find a diffusive scaling $t_{\rm Th}\propto L^{2}$ below the critical disorder strength [^17] In contrast, a sub-diffusive scaling of $t_{Th}$ is found in the delocalized phase of Anderson model on random regular graphs [136].. This scaling is slower than the scaling of $t_{H}$, which in the 3D Anderson model scales as $t_{H}\propto L^{3}$. At the critical point, however, it was found [593, 738] that $t_{\rm Th}\propto t_{H}\propto L^{3}$, and hence the scale invariance (i.e., independence of $L$) of the ratio $t_{H}/t_{\rm Th}$ represents an adequate criterion for the detection of the critical point. In the localized phase, the SFF of the 3D Anderson model tends, with increasing $L$, to the SFF for Poisson level statistics, $K_{PS}(\tau)=1$, and hence the notion of $t_{\rm Th}$ extracted from the SFF becomes less meaningful.
@@
A nontrivial question is whether the scale invariance of $t_{H}/t_{\rm Th}$ also emerges in disordered spin chains such as the $J_{1}$-$J_{2}$ model. Deep in the ergodic regime of the latter model, see Fig. 5(a), $\tau_{\rm Th}$ decreases significantly with the increase of system size $L$, which, given the exponential scaling of $t_{H}$, amounts to approximately a diffusive scaling of the Thouless time, $t_{\rm Th}\propto L^{2}$. On the other hand, $t_{\rm Th}$ increases approximately exponentially with disorder $W$ [692]. As a consequence, at disorder values that are not extremely large (see, e.g., the results in Fig. 5(b) at $W=10$), $t_{\rm Th}$ in finite systems becomes larger than (or comparable to) $t_{H}$. Investigation of the system size $L$ and disorder strength $W$ dependence of the Thouless time in the disordered XXZ and $J_{1}$-$J_{2}$ models led to the following two-parameter scaling hypothesis for the Thouless time [692]
@@
where $t_{0}$ and $\Omega$ are constants. This equation indicates a diffusive scaling of the Thouless time with the diffusion constant, i.e., the inverse of the prefactor in (32), decreasing exponentially with disorder strength $W$. At any finite system size $L$, the Thouless time (32) is upper-bounded by the Heisenberg time $t_{\mathrm{Th}}<t_{\mathrm{H}}=e^{cL}$. Hence, there exists, at any finite $L$, a disorder strength $W_{\mathrm{Th}}^{T}\sim L$ (up to sub-leading logarithmic corrections) at which the scaling (32) breaks down. When $t_{\rm Th}$ becomes close to $t_{H}$, the spectral correlations do not adhere to GOE predictions even at the smallest energy scales, corresponding to time scales of the order of $t_{H}$. Consequently, if $t_{\rm Th}\gtrsim t_{H}$, the short-range level statistics, as captured, e.g., by the average gap ratio $\overline{r}$, deviate from the GOE behavior. Hence, $W_{\mathrm{Th}}^{T}$ defined here is closely related to the disorder strength $W^{T}(L)$ extracted as the point at which the average gap ratio deviates from the GOE value (28).
@@
Nevertheless, the scaling (32) of $t_{\mathrm{Th}}$ with system size $L$ and disorder strength $W$ highlights a fundamental difficulty in understanding of the MBL phase. The Thouless time $t_{\mathrm{Th}}$ increases, to a good approximation, exponentially with the disorder strength $W$, consistently with the dramatic slow down of the many-body dynamics upon the increase of $W$. The slow down of dynamics hinders the differentiation between a very slow approach to thermal equilibrium and ultimately arrested dynamics in the $t,L\to\infty$ limit consistent with the definition of the MBL phase (6). In finite systems, the scaling (32) is not expected to hold when the values of $t_{\rm Th}$ become comparable to those of $t_{H}$ [692]. A subsequent analysis [593, 597] focused on the presence of deviations from the scaling (32), arguing that the diffusive scaling $t_{\rm Th}\propto L^{2}$ may break down at sufficiently large $L$ and $W$. These deviations suggest that the scaling (32), while approximately valid for disordered spin chains, may not be valid in the asymptotic regime $L\to\infty$ at arbitrary large disorder. Hence, (32) per se does not rule out the existence of the MBL phase, similarly to the gap ratio result (28). Moreover, when the diffusive scaling in (32) is replaced by a subdiffusive dependence $t_{\rm Th}\propto L^{z}$ (with $z>2$), the analysis of the results in much less constrained [692], and presently available system sizes do not allow for unambiguous conclusions regarding both the nature of the dynamics at larger values of $W$ as well as the presence of the MBL phase.
In the ergodic phase of generic Floquet models without any local conservation laws, the Thouless time scales logarithmically with system size [336, 109]. This finding is supported numerically by transfer matrix calculations [217, 218], but the logarithmic system size dependence is non-trivial to observe in the exact diagonalization investigations, see [617]. A protocol to measure SFF with the randomized measurements in quantum simulator settings was recently proposed [302].
@@ VI.1.3 Other measures of spectral statistics
A more traditional approach to study quantum chaos manifested by the onset of level repulsion between eigenvalues of complex systems is to consider the level spacing distribution $P(s)$, exemplified by the investigations of heavy nuclei [250], hydrogen atom in magnetic field [166, 212] or quantum-chaotic spin chains [39, 552]. For systems following the predictions of random matrix theory, the level spacing distribution is approximated by the Wigner surmise,
@@
where $A_{\beta}$, $B_{\beta}$ are constants and $\beta=1,2,4$, depending on the symmetry class of the model [181, 182, 183]. For instance, for time reversal symmetric systems, such as (7), $\beta=1$ and $A_{1}=\pi/2$, $B_{1}=\pi/4$. Presence of the level repulsion ensures that $P_{\beta}(s)\stackrel{{\scriptstyle s\to 0}}{{\longrightarrow}}0$. In contrast, the level spacing distribution for Poisson statistics reads
@@
The evolution of the level spacing distribution from (33) to (34) at the crossover between ETH and MBL regimes in the disordered XXZ spin chain has been studied in a number of works [118, 577, 68, 97, 602, 325]. Consistently with the trends visible in the results for the average gap ratio, the level spacing distribution changes significantly with system size $L$ at a fixed disorder strength $W$ [603]. Moreover, to extract the level spacing distribution from eigenvalues of a many-body system, one has to perform a procedure of the so-called unfolding, which fixes the density of energy levels to unity. The unfolding procedure is not uniquely defined and may be performed in several distinct ways [237, 60], which may lead to misleading results. For that reason, the analysis of average gap ratio (26) is simpler and more commonly employed.
@@
We note that the number variance $\Sigma^{2}$, a quantity directly related to two-point correlation function of energy levels [410] was employed in the studies of long-range spectral correlations at the ETH-MBL crossover [68, 602]. The number variance $\Sigma^{2}$, together with the approach proposed in [139], provide alternative routes for the definition of Thouless time based on the properties of long-range spectral statistics.
@@ VI.1.4 Overview
The SFF quantifies two-point long-range spectral correlations, and its characteristic ramp-plateau structure in disordered many-body systems allows to define the Thouless time $t_{\mathrm{Th}}$, beyond which the approach to equilibrium state is over. Investigations of system size and disorder strength scaling of $t_{\mathrm{Th}}$ lead to ambiguous conclusions for the ETH-MBL crossover, similarly to the average gap ratio studies. The approximately exponential increase of $t_{\mathrm{Th}}$ with the disorder strength $W$ reflects the dramatic slow down of the many-body dynamics with increase of the disorder strength. These results, in the regime of $L$ accessible to the present-day numerics, could be consistent both with a slow approach to thermal equilibrium as well as with a complete freezing of the dynamics in the $t,L\to\infty$ limit.
@@ VI.2.1 Wave function properties
The eigenstates |$\ket{m}$ are determined as linear combinations of $\mathcal{N}$ vectors of the many-body basis of the Hilbert space. The eigenbasis |$\ket{\vec{\sigma}}$ of $S^{z}_{i}$ operators, which we refer to also as the computational basis, is a natural basis for the investigation of MBL in the disordered XXZ spin chain (7) or the $J_{1}$-$J_{2}$ model (27). The eigenstates of these disordered spin chains coincide with the basis states |$\ket{\vec{\sigma}}$ in the $W\to\infty$ limit at any fixed system size $L$. Participation entropy, i.e., the Rényi entropy with index $q$ associated with the probability distribution $|\braket{\vec{\sigma}}{m}|^{2}$, defined as
$S_{q}=\frac{1}{1-q}\log_{2}\left(\sum_{\vec{\sigma}}|\braket{\vec{\sigma}}{\psi}|^{2q}\right),$ (36)
allows to study the spread of the state |$\ket{m}$ in the many-body basis |$\ket{\vec{\sigma}}$ at finite $L$ and $W$. Participation entropy is related to the inverse participation ratio, which played a significant role in the analysis of Anderson transitions [655, 344, 197]. Moreover, the participation entropies of many-body ground states distinguish between various quantum phases [627, 625, 626, 392, 396, 487, 376, 476] and have been proposed as an ergodicity breaking measure in [157].
@@
where $D_{q}$ is fractal dimension [263], $c_{q}$ is a sub-leading term, and $\mathcal{N}$ is the dimension of Hilbert space in the relevant symmetry sector, i.e., $\mathcal{N}=\binom{L}{L/2}$ in the $\sum_{i}S^{z}_{i}=0$ sector. In general, (37) may be expected to apply only in a narrow interval of system sizes. When the participation entropy $S_{q}$ is available for a wider interval of system sizes, interpolating the dependence of $S_{q}$ on $\log_{2}(\mathcal{N})$ enables calculation of the fractal dimension as the first derivative $D_{q}=\frac{\partial S_{q}}{\partial(\log_{2}(\mathcal{N}))}$. If the state |$\ket{m}$ is supported on a fixed number of basis states, then $S_{q}$ is independent of $L$ and consequently, the fractal dimension vanishes, $D_{q}=0$. In contrast, a Haar-random state [^18] A Haar-random state can be obtained as |$\ket{\psi}=U\ket{\psi_{0}}$, where |$\ket{\psi_{0}}$ is a fixed state and $U$ is a matrix drawn with Haar measure on a selected matrix group. The orthogonal group, $U\in O(\mathcal{N})$, is relevant for time-reversal invariant systems., which can often be used as an approximation for eigenstates of ergodic many-body systems, is fully delocalized in the Hilbert space, has $D_{q}=1$ (and $c_{q}<0$[41]. Multifractality [621] is the intermediate case when the fractal dimension $10<D_{q}<$ 1 depends non-trivially on the Rényi index $q$.
@@
[…]
@@
The fractal dimension $D_{q}$ exhibits a non-monotonous dependence on $W$, admitting a maximum at disorder strength at the boundary of the ETH regime at a given system size $L$, see Fig. 6. This region of parameters coincides with the regime dubbed as maximally chaotic, in which the sensitivity of the eigenstates to perturbations is maximal [373], see also Sec. VI.3.3. The maximum of $D_{q}$ is followed by its decrease with $W$ towards the value characteristic for the MBL regime. This behavior was interpreted as a transition to MBL phase [401], leading to an estimate for the critical disorder strength consistent with the earlier study [393]. Significantly, however, the crossing point of the curves $D_{q}(L)$ for various system sizes drifts towards larger $W$ with increasing $L$. The drift of the crossing point for $D_{q}$, within the interval of available system sizes, is weaker than the drift for the average gap ratio, c.f. Fig. 4. Nevertheless, this finite size drift of the crossover prevents one from reaching ambiguous conclusions about the fate of the ETH-MBL crossover in the $L\to\infty$ limit.
@@
The multifractal structure of the many-body wave function of eigenstates in the MBL regime can be employed to develop decimation schemes that discard the irrelevant parts of the Hilbert space, facilitating their numerical description [483], see also [499, 504]. The properties of many-body wave function do depend on the choice of the basis. However, using another choice of the eigenbasis, namely, localized Anderson orbitals obtained by diagonalization of (7) in the single particle sector, the results for the $D_{q}$ remain qualitatively the same [401]. Nevertheless, [391] suggests that the dynamical properties of many-body systems do not depend strongly on the degree of multifractality of many-body eigenstates. We note that the participation entropies at the ETH-MBL crossover may not be self-averaging [616] so that extraction of a representative value of $S_{q}$ at disorder strengths in that regime is possible only upon averaging the results over multiple disorder realizations of the system. Similar reservations also apply to other quantities at the ETH-MBL crossover [666].
@@
Non-equilibrium problems for the disordered XXZ spin chain (7) can be viewed as a problem of a fictitious particle hopping on a Fock space graph [379, 532, 641], whose vertices are the spin configurations |$\ket{\vec{\sigma}}$ of $S^{z}_{i}$. The Fock space graph has an extensive local connectivity, and the on-site energies are correlated, which may prevent the wave function of the fictitious particle from trivial delocalization [534, 535]. We note that the effects of Fock space correlations on finite system size drifts of the ETH-MBL crossover in several models was analyzed in [569]. The Fock space perspective enables application of insights and approximations developed for single particle problems to analyze properties of many-body system [536]. Still, the essential properties and system size drifts of the obtained measures are similar [631] to the features typically observed at the crossover, as captured, e.g., by the participation entropy (36).
@@
Figure 7: Entanglement of mid-spectrum eigenstates of the $J_{1}$-$J_{2}$ model at the ETH-MBL crossover. Quantum mutual information $I_{2}$ between pairs of the spins in chain of $L=16$ sites in the ETH regime, (a), and in the MBL regime, (b). The thickness of the line is proportional to the value of $I_{2}$. (c) Rescaled entanglement entropy $\overline{s}$ crossovers between a volume-law scaling $\overline{s}\to 1$ at small disorder strength and an area-law scaling $\overline{s}\to 0$ deep in the MBL regime. Shift of the crossing point (denoted by the red cross) with system size $L$ is significant.

### v3

@@ VI.1.2 Thouless time
Figure 5: Spectral form factor $K(\tau)$ in the $J_{1}$-$J_{2}$ spin-1/2 chain. (a) Deep in the ergodic regime, for $W=2.5$, the SFF of the $J_{1}$-$J_{2}$ model has the ramp-plateau structure following the GOE prediction (31) (denoted by gray dashed line) for $\tau>t_{Th}/t_{H}$, defining the Thouless time $t_{Th}$. The points $\tau_{Th}=t_{Th}/t_{H}$ are denoted by red circles, while the red square denotes $\tau_{\mathrm{H}}=1$. (b) Deep in the MBL regime of $J_{1}$-$J_{2}$ model, for $W=10$, the SFF is slowly tending, with an increase of $L$, to the result $K_{PS}(\tau)=1$ for Poisson level statistics. Convergence of SFF to $K_{PS}(\tau)$ for $L\to\infty$ at $W>W_{c}$ would imply the presence of the MBL phase in the system.
@@
The spectral form factor (SFF) is a Fourier transform of a two-point spectral correlation function. It quantifies long-range correlations between eigenvalues of many-body systems, and it provides insight into the dynamical properties at timescales below the Heisenberg time $t_{H}$. Investigations of the SFF played a significant role in our understanding of the ETH-MBL crossover. If we denote the unfolded [^24] Spectral unfolding is a procedure of setting the density of eigenvalues to unity [274, 253], see also Sec. VI.1.3. eigenvalues of the considered quantum system by $\{\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{\mathcal{N}}\}$, the averaged SFF is defined as
@@
where the average is taken over disorder realizations. The filter function $g(\epsilon)$ in Eq. (30) is a Gaussian function with the mean in the middle of the unfolded spectrum and with the variance proportional to the variance of $\{\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{\mathcal{N}}\}$, with the proportionality factor $\frac{3}{10}$, following the choice of [746]. The filter $g(\epsilon)$ reduces the influence of the spectrum’s edges, while $Z$ is a normalization factor ensuring that $K(\tau)\stackrel{{\scriptstyle\tau\to\infty}}{{\longrightarrow}}1$. The SFF is known analytically for the Gaussian ensembles of random matrices. In particular, for the GOE, it reads [446]:
@@
[…]
@@
A nontrivial question is whether the scale invariance of $t_{H}/t_{\rm Th}$ also emerges in disordered spin chains such as the $J_{1}$-$J_{2}$ model. Deep in the ergodic regime of the latter model, see Fig. 5(a), $\tau_{\rm Th}$ decreases significantly with the increase of system size $L$, which, given the exponential scaling of $t_{H}$, amounts to approximately a diffusive scaling of the Thouless time, $t_{\rm Th}\propto L^{2}$. On the other hand, $t_{\rm Th}$ increases approximately exponentially with disorder $W$ [748]. As a consequence, at disorder values that are not extremely large (see, e.g., the results in Fig. 5(b) at $W=10$), $t_{\rm Th}$ in finite systems becomes larger than (or comparable to) $t_{H}$. Investigation of the system size $L$ and disorder strength $W$ dependence of the Thouless time in the disordered XXZ and $J_{1}$-$J_{2}$ models led to the following two-parameter scaling hypothesis for the Thouless time [748],
@@
where $t_{0}$ and $\Omega$ are constants. This equation indicates a diffusive scaling of the Thouless time, with the diffusion constant (that can be interpreted as the inverse of the prefactor in (32)) decreasing exponentially with disorder strength $W$ [^27] This scaling is inconsistent with the recent mathematical arguments of [573] for the disordered quantum Ising chain, which rule out diffusive transport at sufficiently large disorder strengths $W$. . At any finite system size $L$, the Thouless time (32) is upper-bounded by the Heisenberg time $t_{\mathrm{Th}}<t_{\mathrm{H}}=e^{cL}$. Hence, there exists, at any finite $L$, a disorder strength $W_{\mathrm{Th}}^{T}\sim L$ (up to sub-leading logarithmic corrections) at which the scaling (32) breaks down. When $t_{\rm Th}$ becomes close to $t_{H}$, the spectral correlations do not adhere to GOE predictions even at the smallest energy scales, corresponding to time scales of the order of $t_{H}$. Consequently, if $t_{\rm Th}\gtrsim t_{H}$, the short-range level statistics, as captured, e.g., by the average gap ratio $\overline{r}$, deviate from the GOE behavior. Hence, $W_{\mathrm{Th}}^{T}$ defined here is closely related to the disorder strength $W^{T}(L)$ extracted as the point at which the average gap ratio deviates from the GOE value (28).
@@
Nevertheless, the scaling (32) of $t_{\mathrm{Th}}$ with system size $L$ and disorder strength $W$ highlights a fundamental difficulty in understanding of the MBL phase. The Thouless time $t_{\mathrm{Th}}$ increases, to a good approximation, exponentially with the disorder strength $W$, consistently with the dramatic slow down of the many-body dynamics upon the increase of $W$. The slow down of the dynamics hinders the differentiation between a very slow approach to thermal equilibrium and ultimately arrested dynamics in the $t,L\to\infty$ limit, consistent with the definition of the MBL phase (6). In finite systems, the scaling (32) is not expected to hold when the values of $t_{\rm Th}$ become comparable to those of $t_{H}$ [748]. A subsequent analysis [642, 646] focused on the presence of deviations from the scaling (32), arguing that the diffusive scaling $t_{\rm Th}\propto L^{2}$ may break down at sufficiently large $L$ and $W$. These observations suggest that the scaling (32), while approximately valid for disordered spin chains, may not be valid in the asymptotic regime $L\to\infty$ at arbitrary large disorder. Hence, (32) per se does not rule out the existence of the MBL phase, similarly to the gap ratio result (28). Moreover, when the diffusive scaling in (32) is replaced by a subdiffusive dependence $t_{\rm Th}\propto L^{z}$ (with $z>2$), the analysis of the results in much less constrained [748], and presently available system sizes do not allow for unambiguous conclusions regarding both the nature of the dynamics at larger values of $W$ as well as the presence of the MBL phase.
In the ergodic phase of generic Floquet models without any local conservation laws, the Thouless time scales logarithmically with system size [364, 113]. This finding is supported numerically by transfer matrix calculations [232, 233], but the logarithmic system size dependence is non-trivial to observe in the exact diagonalization investigations, see [667]. A protocol to measure SFF with the randomized measurements in quantum simulator settings was recently proposed [328] and executed for small quantum systems in [185].
@@ VI.1.3 Other measures of spectral statistics
A more traditional approach to study quantum chaos, manifested by the onset of level repulsion between eigenvalues of complex systems, is to consider the level spacing distribution $P(s)$, exemplified by the investigations of heavy nuclei [266], hydrogen atom in magnetic field [174, 226] or quantum-chaotic spin chains [41, 598]. For systems following the predictions of random matrix theory, the level spacing distribution is approximated by the Wigner surmise,
@@
where $A_{\beta}$, $B_{\beta}$ are constants and $\beta=1,2,4$, depending on the symmetry class of the model [191, 192, 193]. For instance, for time reversal symmetric systems, such as (7), $\beta=1$ and $A_{1}=\pi/2$, $B_{1}=\pi/4$. Emergence of the level repulsion ensures that $P_{\beta}(s)\stackrel{{\scriptstyle s\to 0}}{{\longrightarrow}}0$. In contrast, the level spacing distribution for Poisson statistics reads
@@
The evolution of the level spacing distribution from (33) to (34) at the crossover between ETH and MBL regimes in the disordered XXZ spin-1/2 chain has been studied in a number of works [124, 624, 71, 100, 651, 352]. Consistently with the trends visible in the results for the average gap ratio, the level spacing distribution changes significantly with system size $L$ at a fixed disorder strength $W$ [652]. Moreover, to extract the level spacing distribution from eigenvalues of a many-body system, one has to perform a procedure of the so-called unfolding, which fixes the density of energy levels to unity. The unfolding procedure is not uniquely defined and may be performed in several distinct ways [253, 63], which may lead to misleading results, i.e., the probes of quantum chaotic behavior, especially related to long-range spectral statistics, may strongly depend on the employed unfolding procedure [253]. For that reason, the analysis of the average gap ratio (26) is simpler and more commonly employed.
@@
We note that the number variance $\Sigma^{2}$, a quantity directly related to two-point correlation function of energy levels [446], was employed in the studies of long-range spectral correlations at the ETH-MBL crossover [71, 651]. The number variance $\Sigma^{2}$, together with the approach proposed in [144], provide alternative routes for the definition of Thouless time based on the properties of long-range spectral statistics.
@@ VI.1.4 Overview
The SFF quantifies two-point long-range spectral correlations, and its characteristic ramp-plateau structure in disordered many-body systems allows to define the Thouless time $t_{\mathrm{Th}}$, beyond which the approach to equilibrium state is over. Investigations of system size and disorder strength scaling of $t_{\mathrm{Th}}$ lead to ambiguous conclusions for the ETH-MBL crossover, similarly to the average gap ratio studies. The approximately exponential increase of $t_{\mathrm{Th}}$ with the disorder strength $W$ reflects the dramatic slow down of the many-body dynamics with increase of the disorder strength. These results, in the regime of system sizes $L$ accessible to the present-day numerics, could be consistent both with a slow approach to thermal equilibrium as well as with a complete freezing of the dynamics in the $t,L\to\infty$ limit.
@@ VI.2.1 Wave function properties
The eigenstates |$\ket{m}$ are determined as linear combinations of $\mathcal{N}$ vectors of the many-body basis of the Hilbert space. The eigenbasis |$\ket{\vec{\sigma}}$ of $\hat{S}^{z}_{i}$ operators, which we refer to also as the computational basis, is a natural basis for the investigation of MBL in the disordered XXZ spin chain (7) or the $J_{1}$-$J_{2}$ model (27). The eigenstates of these disordered spin chains coincide with the basis states |$\ket{\vec{\sigma}}$ in the $W\to\infty$ limit at any fixed system size $L$. For finite $W$, the participation entropy, i.e., the Rényi entropy with index $q$ associated with the probability distribution $|\braket{\vec{\sigma}}{m}|^{2}$, defined as
$S_{q}=\frac{1}{1-q}\log_{2}\left(\sum_{\vec{\sigma}}|\braket{\vec{\sigma}}{m}|^{2q}\right),$ (36)
allows to study the spread of the state |$\ket{m}$ in the many-body basis |$\ket{\vec{\sigma}}$ which, for $U(1)$ symmetric models, spans a subspace with fixed magnetization (in the following, we focus on the zero magnetization sector). Participation entropy is related to the inverse participation ratio, which played a significant role in the analysis of Anderson transitions [707, 373, 207]. Moreover, the participation entropies of many-body ground states distinguish between various quantum phases [677, 675, 676, 429, 432, 526, 411, 514] and have been proposed as an ergodicity breaking measure in [164]. Quantum algorithms for measuring the participation entropies were proposed by [412].
@@
where $D_{q}$ is fractal dimension [279], $c_{q}$ is a sub-leading term, and $\mathcal{N}$ is the dimension of Hilbert space in the relevant symmetry sector, i.e., $\mathcal{N}=\binom{L}{L/2}$ in the $\sum_{i}\hat{S}^{z}_{i}=0$ sector. In general, (37) may be expected to apply only in a narrow interval of system sizes. When the participation entropy $S_{q}$ is available for a wider interval of system sizes, interpolating the dependence of $S_{q}$ on $\log_{2}(\mathcal{N})$ enables calculation of the fractal dimension as the first derivative, $D_{q}=\frac{\partial S_{q}}{\partial(\log_{2}(\mathcal{N}))}$. If the state |$\ket{m}$ is supported on a fixed number of basis states, then $S_{q}$ is independent of $L$ and consequently, the fractal dimension vanishes, $D_{q}=0$. In contrast, a Haar-random state [^28] A Haar-random state can be obtained as |$\ket{\psi}=U\ket{\psi_{0}}$, where |$\ket{\psi_{0}}$ is a fixed state and $U$ is a matrix drawn with Haar measure on a selected matrix group. The orthogonal group, $U\in O(\mathcal{N})$, is relevant for time-reversal invariant systems., which can often be used as an approximation of eigenstates of an ergodic many-body system, if fully delocalized in the Hilbert space, has $D_{q}=1$ (and $c_{q}<0$[43]. Multifractality [671] is the intermediate case when the fractal dimension $10<D_{q}<$ 1 depends non-trivially on the Rényi index $q$.
@@
For the $J_{1}$-$J_{2}$ model, at system size $L$, the fractal dimensions $D_{q}$ delineate a pronounced ETH regime, in which $D_{q}=1$, and the MBL regime, in which the states are multifractal and $10<D_{q}<$ 1, depending non-trivially on the Rényi index $q$, and decreasing approximately as $D_{q}\propto 1/W$ at $W\gg 1$ [437]. Notably, this implies that the eigenstates in the MBL regime are not localized in the computational basis |$\ket{\vec{\sigma}}$, as already anticipated in [98, 163]. Considering many-body eigenstates in the Anderson insulator, by setting $\Delta=0$ in (7), one finds that even in the absence of interactions, the states |$\ket{m}$ are not localized in the |$\ket{\vec{\sigma}}$ basis (in the sense of vanishing $D_{q}$), but rather are multifractal [^29] To the best of our knowledge, a demonstration of this fact in large systems of free fermions is presently lacking.. From this perspective, it is natural to expect that the eigenstates are still multifractal (in the computational basis) in the MBL regime upon the introduction of interactions to the Anderson insulator. For this reason, if an MBL phase exists beyond a certain critical disorder strength $W_{c}$ in the considered spin chains, it is natural to expect that it is associated with multifractal eigenstates in the computational basis. In contrast, localization of the states in computational basis |$\ket{\vec{\sigma}}$, denoted also as the Fock-space localization, occurring when $D_{q}=0$ 14, is a stronger form of ergodicity breaking, which may arise in disordered spin chains at disorder strengths larger than $W\propto L$ [171], and occurs in the quantum sun model [744].
@@
The fractal dimension $D_{q}$ exhibits a non-monotonous dependence on $W$, admitting a maximum at disorder strength at the boundary of the ETH regime at a given system size $L$, see Fig. 6. The fractal dimension $D_{q}$ can be bigger than unity since it reflects the local slope of the $S_{q}(L)$ curve. The interval of disorder strengths $W$, for which $D_{q}>1$, shrinks with the increase of $L$. This region of parameters coincides with the regime dubbed as maximally chaotic, in which the sensitivity of the eigenstates to perturbations is maximal [408], see also Sec. VI.3.3. The maximum of $D_{q}$ is followed by its decrease with $W$ towards the value characteristic for the MBL regime. This behavior was interpreted as a transition to MBL phase [437], leading to an estimate for the critical disorder strength consistent with the earlier study [430]. Significantly, however, the crossing point of the curves $D_{q}(L)$ for various system sizes drifts towards larger $W$ with increasing $L$. The drift of the crossing point for $D_{q}$, within the interval of available system sizes, is weaker than the drift for the average gap ratio, c.f. Fig. 4. Nevertheless, this finite size drift of the crossover prevents one from reaching unambiguous conclusions about the fate of the ETH-MBL crossover in the $L\to\infty$ limit.
@@
The multifractal structure of the many-body wave function of eigenstates in the MBL regime can be employed to develop decimation schemes that discard the irrelevant parts of the Hilbert space, facilitating their numerical description [522], see also [540, 545]. The properties of many-body wave function do depend on the choice of the basis. However, using another choice of the eigenbasis, namely, localized Anderson orbitals obtained by diagonalization of (7) in the single particle sector, the results for $D_{q}$ were reported to remain qualitatively the same [437]. Nevertheless, [428] suggests that the dynamical properties of many-body systems do not depend strongly on the degree of multifractality of many-body eigenstates. We note that the participation entropies at the ETH-MBL crossover may not be self-averaging [666], so that the extraction of a representative value of $S_{q}$ at disorder strengths in that regime is possible only upon averaging the results over multiple disorder realizations of the system. Similar reservations also apply to other quantities at the ETH-MBL crossover [719].
@@
Non-equilibrium problems for the disordered XXZ spin chain (7) can be viewed as a problem of a fictitious particle hopping on a Fock space graph [415, 578, 692], whose vertices are the spin configurations |$\ket{\vec{\sigma}}$ corresponding to eigenstates of the $\hat{S}^{z}_{i}$ operators. The Fock space graph has an extensive local connectivity, and the on-site energies are correlated, which may prevent the wave function of the fictitious particle from trivial delocalization [580, 581]. We note that the effects of Fock space correlations on finite system size drifts of the ETH-MBL crossover in several models was analyzed in [616]. The Fock space perspective enables application of insights and approximations developed for single particle problems to analyze properties of many-body system [582]. Still, the essential properties and system size drifts of the obtained measures are similar [681] to the features typically observed at the crossover, as captured, e.g., by the participation entropy (36).
@@
Figure 7: Entanglement of mid-spectrum eigenstates of the $J_{1}$-$J_{2}$ model at the ETH-MBL crossover. Quantum mutual information $I_{2}$ between pairs of the spins in chain of $L=16$ sites in the ETH regime, (a), and in the MBL regime, (b). The thickness of the line is proportional to the value of $I_{2}$. Rescaled entanglement entropy $\overline{s}$ crossovers between a volume-law scaling $\overline{s}\to 1$ at small disorder strength and an area-law scaling $\overline{s}\to 0$ deep in the MBL regime. Shift of the crossing point (denoted by the red cross) with system size $L$ is significant.